OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit. An=n/(1+ sqrt(n))

OpenStudy (anonymous):

use nth term test (also called Divergence Test) if \[\Large \lim_{n \rightarrow \infty} a_{n} \ne 0\] then series is divergent so \[\Large \lim_{n \rightarrow \infty} \frac{n}{1+\sqrt{n}}\] can you find this Limit ?

OpenStudy (anonymous):

it goes to infinite right?

OpenStudy (anonymous):

yes correct . so it means the series is .............. ?

OpenStudy (anonymous):

divergent

OpenStudy (anonymous):

yes thats correct !

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

:) you are welcome !